OpenStudy (anonymous):

Evaluate the difference quotient for the given function. Simplify your answer. f(x) = x + 8 x + 2 , f(x) − f(4) x − 4

OpenStudy (anonymous):

difference quotient: \[\frac{ f \left( x+h \right)-f \left( x \right) }{ h }\]is typically like this. but for yours, simply do this: \[\frac{ f \left( x \right)-f \left( 4 \right) }{ x-4 }\]

OpenStudy (anonymous):

is it f(x) = x + 8x + 2 = 9x + 2 or is the first term supposed to be \(x^2\)?

OpenStudy (anonymous):

I get as far as \[(\frac{ x+8 }{x+2 }-2)\div(x-4)\]

OpenStudy (anonymous):

what is the function?

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

\[\frac{ x+8 }{ x+2 }\] is the function

OpenStudy (anonymous):

okay, but write it like this:\[\frac{ \frac{ x+8 }{ x+2 }-2 }{ x-4 }\]then simplify the top and then divide by (x-4). remember that you can multiply by the reciprocal to do the division.